• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 00:21:45 on December24,2003

    Entry number 115132

    keyword=Shift Summary

    Beam steering test of left septum shows very little steering of the primary beam with 187 amps in the left septum.
    Then we did RICH studies which are summarized in Bodo's halogs

    Found that the RICH HV tripped very consistently when the elastic peak was located at about delta = +1% (either the edge of the middle cathode or the edge of the next cathode) and continued for all higher values of delta.

    Meanwhile, Bogdan does his own independant detector studies. ("A1 is not in bad shape" "A2 unfortunately the same as 2 years ago")